if a+b+c =0 then what is the value of a^2+b^2+c^2 / c^2-ab 

Grade:12

1 Answers

HONEY
5 years ago
c²=a²+b²+ab as c=-a-b
Now a²+b²+c² = -2(ab+bc+ac)
So -2 (ab +c(a+b)) and c=-a-b 
Now putting value of c and multiplying we get
2(a²+b²+ab) /(a²+b²+ab) which gives answer equal to 2.